OpenStudy (anonymous):

a child of mass 50.0kg sitting 1.0m from the center of a merry-go-round moves with a speed of 1.5 m/s. calculate the centripetal acceleration of the child

OpenStudy (anonymous):

We just need to use a common equation involved with circular motion. The main equation we should know is:\[a_{c}=\frac{ v^{2} }{ r }\]We are given m, v, and r in the problem. However, notice that centripetal acceleration is not dependent on the mass of the object!\[a_{c}=\frac{ v^{2} }{ r }=\frac{ (1.5)^{2} }{ 1.0 }=1.5^{2} =2.25m/s^{2}\]The mass of the child is extra info. Unless you are then asked to find the centripetal force, you would just use:\[F_{c}=ma_{c}\]